OpenStudy (anonymous):

Somebody please help, I'm clueless. Write the equation for a translation to the right 2 units and then down by 1 unit to f(x)=x^3+2x^2+x-3

OpenStudy (anonymous):

oops right 2 units, ok us \[f(x-2)\]

OpenStudy (anonymous):

then down one, find \[f(x-2)-1\]

OpenStudy (anonymous):

I don't get it?

OpenStudy (lastdaywork):

What satellite73 means is - new function - f(x)=[(x-2)^3+2(x-2)^2+(x-2)-3] - 1

OpenStudy (anonymous):

So that's the answer?

OpenStudy (lastdaywork):

Yes, although you still have to simplify it.

OpenStudy (anonymous):

Well I know how to simplify, but I don't understand how you got the new function.

OpenStudy (lastdaywork):

Let there be a graph y=f(x) Horizontal transformation ('a' units towards +ive x-axis) - y=f(x-a) Vertical transformation ('a' units towards +ive y-axis) - (y-a)=f(x) although most people write it as - y=f(x) + a There are many more transformations, I can cite some books if you want (completely optional :P )

OpenStudy (lastdaywork):

@TheJonathanP4 So what is it that you still don't understand ?

OpenStudy (anonymous):

I'm just not positive what I exactly write down for the answer. Do I write f(x)=[(x-2)^3+2(x-2)^2+(x-2)-3] - 1

OpenStudy (anonymous):

@LastDayWork

OpenStudy (lastdaywork):

You have to simplify the expression in RHS. That will be your final answer.

OpenStudy (anonymous):

what is RHS

OpenStudy (lastdaywork):

Shorthand for - Right Hand Side of the equation.

OpenStudy (anonymous):

How do I do that? do you mean change it to look like this f(x-2)-1=x^3+2x^2+x-3

OpenStudy (anonymous):

@phi Can you please help explain what they're asking me to do.

OpenStudy (phi):

We can show by pictures.

OpenStudy (phi):

the three curves are: black (the original curve f(x)) red = the black curve shifted down. we add -1 to f(x) green = black curve shifted up. we add +3 to f(x)

OpenStudy (phi):

Does that make sense? to shift a curve up U steps, add +U to the curve. All the y values become U bigger. The x values stay the same.

OpenStudy (anonymous):

yes, I understand that

OpenStudy (phi):

so do *part of the problem* what is f(x)=x^3+2x^2+x-3 shifted *down* by one step ?

OpenStudy (phi):

shifted down means you want f(x) to be smaller by 1 step. In other words you want f(x) - 1 that means write down f(x) f(x)=x^3+2x^2+x-3 and add -1 to both sides f(x) - 1 = x^3+2x^2+x-3 - 1 or f(x) - 1 = x^3+2x^2+x-4

OpenStudy (anonymous):

okay I see

OpenStudy (phi):

what do you get if you shift f(x) up by 2 ?

OpenStudy (anonymous):

to shift up add outside the function f(x)+b; 'b' being the number of units shifted up. so add 2 to the outside

OpenStudy (phi):

ok. now what about shifting sideways ?

OpenStudy (anonymous):

to shift left add inside f(x+b) to shift right f(x-b)

OpenStudy (anonymous):

so since i have to move to the right 2 units and down 1 unit it would look like f(x-2)-1=x^3+2x^2+x-3

OpenStudy (phi):

You also have to replace "x" in the original right hand side with "(x-2)"

OpenStudy (anonymous):

so f(x-2)-1=(x-2)^3+2(x-2)^2+(x-2)-3

OpenStudy (anonymous):

@phi

OpenStudy (phi):

yes, but you are missing a -1 at the very end. you start with f(x)=x^3+2x^2+x-3 shift down 1: f(x) - 1 = x^3+2x^2+x-3 - 1 now shift to the right 2 steps: f(x-2)-1 = (x-2)^3 + 2(x-2)^2 +(x-2) -3 -1 of course you can simplify that. Here is what both curves look like

OpenStudy (phi):

it takes a lot of work, but f(x-2)-1 = (x-2)^3 + 2(x-2)^2 +(x-2) -3 -1 can be re-written as f(x-2)-1 = x^3-4x^2+5x - 6

OpenStudy (anonymous):

what is the identity that will prove 25-4=21